### 06 Mar 2026 v1

@@ 6 Conclusion
(29) $2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{F}(\mathbf{x}_{i}))\|_{2}^{2}\right]\leq 2n\varepsilon^{2}~.$
@@
(30) $\chi_{1}\leq 4D\varepsilon^{2}+2n\varepsilon^{2}~.$
@@
$T_{1}:=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{G}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\bigr\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\bigr\|_{2}^{2}\right].$
According to [25, Lemma 2], for any $0<\delta<1$ we have
(31) $T_{1}\leq\frac{35DB^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{2DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)+6\delta,$
where $\|f\|_{L^{\infty,\infty}}:=\sup\limits_{\mathbf{x}}\|f(\mathbf{x})\|_{\infty}$.
@@
$\leq$ $\sup_{x\in\mathcal{M}}\Big\|\mathscr{D}_{1}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(x)-\mathscr{D}_{2}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(x)\Big\|_{\infty}$
$+\sup_{x\in\mathcal{M}}\Big\|\mathscr{D}_{2}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(x)-\mathscr{D}_{2}\circ\mathbf{A}_{2}\circ\mathscr{E}_{2}(x)\Big\|_{\infty}$
$\leq$ $\|\mathscr{D}_{1}-\mathscr{D}_{2}\|_{L^{\infty,\infty}}+C_{\phi}\sup_{x\in\mathcal{M}}\Big\|\mathbf{A}_{1}\circ\mathscr{E}_{1}(x)-\mathbf{A}_{2}\circ\mathscr{E}_{2}(x)\Big\|_{2}$
$\leq$ $\|\mathscr{D}_{1}-\mathscr{D}_{2}\|_{L^{\infty,\infty}}+C_{\phi}\sup_{x\in\mathcal{M}}\Big(\|\mathbf{A}_{1}\|_{2}\,\|\mathscr{E}_{1}(x)-\mathscr{E}_{2}(x)\|_{2}+\|\mathbf{A}_{1}-\mathbf{A}_{2}\|_{2}\,\|\mathscr{E}_{2}(x)\|_{2}\Big)$
@@
$\log\mathcal{Q}\!\left(\frac{\delta}{2DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)\leq\;$ $\mathcal{O}\!\left(L_{\mathscr{E}}K_{\mathscr{E}}\bigl(\log p_{\mathscr{E}}+\log\kappa_{\mathscr{E}}+\log\delta_{\mathscr{E}}^{-1}\bigr)\right)$
@@
Lemma  A.3 also satisfies the architecture scaling in [25, Lemma 6 and Lemma 8], we finally obtain
$\log\mathcal{Q}\!\left(\frac{\delta}{2DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)$
@@
$T_{2}:=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\bigr\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\bigr\|_{2}^{2}\right],$
and
$T_{3}:=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}(\mathbf{x})\bigr\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{x}_{i})\bigr\|_{2}^{2}\right].$
Similarly, we obtain that for any $0<\delta<1$,
(34) $T_{2}\leq\frac{35DB^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{2DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{DE}},\|\cdot\|_{L^{\infty,\infty}}\right)+6\delta\leq\mathcal{O}\!\left(D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N\right),$
(35) $T_{3}\leq\frac{35n\Lambda^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{2n\Lambda},\mathcal{F}_{\mathrm{AE}}^{\mathscr{H}},\|\cdot\|_{L^{\infty,\infty}}\right)+6\delta\leq\mathcal{O}\!\left(nD(\log^{2}D)N^{-\frac{2}{n+2}}\log^{4}N\right).$
@@
$\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\Bigl[\|\widehat{\mathscr{G}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\|_{2}^{2}+\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\|_{2}^{2}+\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}(\mathbf{x})\|_{2}^{2}\Bigr]$
$\leq\;$ $4DN^{-\frac{2}{n+2}}+2nN^{-\frac{2}{n+2}}+C_{2}\,D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N$
(36) $\leq\;$ $C\,D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N,$
@@
for some constant $C$ depending on $n,C_{\varphi},C_{\phi},\Lambda,B,\rho$, and the volume of $\mathcal{M}$, which completes the proof.

### 29 Apr 2026 v2

@@ 6 Conclusion
(25) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}_{i}^{+}}\right)\|_{2}^{2}\right]\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{2}\varepsilon^{2}+2C_{1}^{2}C_{\varphi}^{2}\mathbb{E}\|\mathbf{w}\|_{2}^{2}~,}$
where constant $C_{2}=C_{1}^{2}\left(\frac{1}{C_{1}-1}+\frac{8}{(\rho+1)^{2}}\right)$.
@@
(26) $\chi_{1}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2C_{1}D+C_{2})\varepsilon^{2}+C_{1}(1+2C_{1}C_{\varphi}^{2})\mathbb{E}\|\mathbf{w}\|_{2}^{2}}~.$
@@
$T_{1}:=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{G}}(\mathbf{x})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}^{+}}\bigr\|_{2}^{2}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}_{i}^{+}}\bigr\|_{2}^{2}\right].$
According to [31, Lemma 2], the result can be extended from the case $C_{1}=2$ to any $C_{1}>1$ by a straightforward modification of the proof. In particular, replacing the constant $\frac{1}{8DB^{2}}$ in (66) therein with $\frac{C_{1}-1}{4C_{1}DB^{2}}$ yields that for any $0<\delta<1$, there exists a constant $C_{3}=\frac{35C_{1}}{2}$ such that
(27) $T_{1}\leq\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{3}}DB^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3C_{1}}\delta.$
@@
$\leq$ $\,\sup_{\mathbf{x}\in\mathcal{M}}\Big\|\mathscr{D}_{1}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(\mathbf{x})-\mathscr{D}_{2}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(\mathbf{x})\Big\|_{\infty}$
$\,+\sup_{\mathbf{x}\in\mathcal{M}}\Big\|\mathscr{D}_{2}\circ\mathbf{A}_{1}\circ\mathscr{E}_{1}(\mathbf{x})-\mathscr{D}_{2}\circ\mathbf{A}_{2}\circ\mathscr{E}_{2}(\mathbf{x})\Big\|_{\infty}$
$\leq$ $\;\|\mathscr{D}_{1}-\mathscr{D}_{2}\|_{L^{\infty,\infty}}+C_{\phi}\sup_{\mathbf{x}\in\mathcal{M}}\Big\|\mathbf{A}_{1}\circ\mathscr{E}_{1}(\mathbf{x})-\mathbf{A}_{2}\circ\mathscr{E}_{2}(\mathbf{x})\Big\|_{2}$
$\leq$ $\;\|\mathscr{D}_{1}-\mathscr{D}_{2}\|_{L^{\infty,\infty}}+C_{\phi}\sup_{x\in\mathcal{M}}\Big(\|\mathbf{A}_{1}\|_{2}\,\|\mathscr{E}_{1}(\mathbf{x})-\mathscr{E}_{2}(\mathbf{x})\|_{2}+\|\mathbf{A}_{1}-\mathbf{A}_{2}\|_{2}\,\|\mathscr{E}_{2}(\mathbf{x})\|_{2}\Big)$
@@
$\log\mathcal{Q}\!\left(\frac{\delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)\leq\;$ $\mathcal{O}\!\left(L_{\mathscr{E}}K_{\mathscr{E}}\bigl(\log p_{\mathscr{E}}+\log\kappa_{\mathscr{E}}+\log\delta_{\mathscr{E}}^{-1}\bigr)\right)$
@@
Lemma  A.3 also satisfies the architecture scaling in [31, Lemma 6 and Lemma 8], we finally obtain
$\log\mathcal{Q}\!\left(\frac{\delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{G}},\|\cdot\|_{L^{\infty,\infty}}\right)$
@@
$T_{2}:=$ $\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\bigr\|_{2}^{2}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\bigr\|_{2}^{2}\right],$
$T_{3}:=$ $\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\bigl\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}^{+}})\bigr\|_{2}^{2}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\mathbb{E}_{\mathcal{S}}\!\left[\frac{1}{N}\sum_{i=1}^{N}\bigl\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}^{+}_{i}})\bigr\|_{2}^{2}\right].$
Similarly, for any $0<\delta<1$, we obtain the following
$T_{2}\leq$ $\,\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{3}}DB^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}DB},\mathcal{F}_{\mathrm{NN}}^{\mathscr{DE}},\|\cdot\|_{L^{\infty,\infty}}\right)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3C_{1}}\delta$
(30) $\leq$ $\,\mathcal{O}\!\left(D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N\right),$
$T_{3}\leq$ $\,\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{3}}n\Lambda^{2}}{N}\log\mathcal{Q}\!\left(\frac{\delta}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}n\Lambda},\mathcal{F}_{\mathrm{AE}}^{\mathscr{H}},\|\cdot\|_{L^{\infty,\infty}}\right)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3C_{1}}\delta$
(31) $\leq$ $\,\mathcal{O}\!\left(nD(\log^{2}D)N^{-\frac{2}{n+2}}\log^{4}N\right).$
@@
$\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\Bigl[\|\widehat{\mathscr{G}}(\mathbf{x})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}^{+}}\|_{2}^{2}+\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\|_{2}^{2}+\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}^{+}})\|_{2}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+R(\widehat{\mathbf{A}})}\Bigr$]
$\leq\;$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2C_{1}D+C_{2})N^{-\frac{2}{n+2}}+C_{1}(1+2C_{1}C_{\varphi}^{2})\mathbb{E}\|\mathbf{w}\|_{2}^{2}}+C_{0}\,D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N$
$\leq\;$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}(1+2C_{1}C_{\varphi}^{2})\mathbb{E}\|\mathbf{w}\|_{2}^{2}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{0}}\,D^{2}(\log^{2}D)\,N^{-\frac{2}{n+2}}\log^{4}N,$
@@
for some constant $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{0}$} depending on $n,C_{1},C_{\varphi},C_{\phi},\Lambda,B,\rho$, and the volume of $\mathcal{M}$, which completes the proof.
Appendix B Proof of Theorem 4.2
Proof B.1.
Let $\mathcal{F}_{\mathrm{NN}}^{obs}(D_{y},m;L_{obs},p_{obs},K_{obs},\kappa_{obs},R_{obs})$ denote the observation encoder class. We follow the same decomposition as in Theorem 4.1:
$\chi_{1}^{obs}$ $=C_{1}\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{E}}_{obs}(\mathbf{y}_{i})-\mathbf{H}\,\mathscr{E}(\mathbf{x}_{i})\|_{2}^{2}\right],$
$\chi_{2}^{obs}$ $=\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{y})}\|\widehat{\mathscr{E}}_{obs}(\mathbf{y})-\mathbf{H}\,\mathscr{E}(\mathbf{x})\|_{2}^{2}-C_{1}\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{E}}_{obs}(\mathbf{y}_{i})-\mathbf{H}\,\mathscr{E}(\mathbf{x}_{i})\|_{2}^{2}\right].$
By optimality of $\widehat{\mathscr{E}}_{obs}$,
$\chi_{1}^{obs}\leq C_{1}\inf_{\mathscr{E}_{obs}\in\mathcal{F}_{\mathrm{NN}}^{obs}}\mathbb{E}_{(\mathbf{x},\mathbf{y})}\|\mathscr{E}_{obs}(\mathbf{y})-\mathbf{H}\mathscr{E}(\mathbf{x})\|_{2}^{2}.$
By the approximation assumption on $\mathcal{F}_{\mathrm{NN}}^{obs}$, there exists $\widetilde{\mathscr{E}}_{obs}\in\mathcal{F}_{\mathrm{NN}}^{obs}$ such that
$\mathbb{E}_{(\mathbf{x},\mathbf{y})}\|\widetilde{\mathscr{E}}_{obs}(\mathbf{y})-\mathbb{E}[\mathbf{H}\mathscr{E}(\mathbf{x})\mid\mathbf{y}]\|_{2}^{2}\leq\varepsilon^{2}.$
Using the standard bias–variance decomposition
$\|\widetilde{\mathscr{E}}_{obs}(\mathbf{y})-\mathbf{H}\,\mathscr{E}(\mathbf{x})\|_{2}^{2}\leq\;2\|\widetilde{\mathscr{E}}_{obs}(\mathbf{y})-\mathbb{E}[\mathbf{H}\,\mathscr{E}(\mathbf{x})\mid\mathbf{y}]\|_{2}^{2}+2\|\mathbb{E}[\mathbf{H}\,\mathscr{E}(\mathbf{x})\mid\mathbf{y}]-\mathbf{H}\,\mathscr{E}(\mathbf{x})\|_{2}^{2}.$
Under a stability assumption on the observation operator $\mathcal{H}$, the conditional variance term can be controlled by the observation noise level. Thus, there exists a constant $C_{1}^{obs}=2C_{1}\|\mathbf{H}\|_{2}^{2}C_{\phi}^{2}C_{\mathcal{H}}^{2}$, such that
$\chi_{1}^{obs}\leq 2C_{1}\varepsilon^{2}+C_{1}^{obs}\,\mathbb{E}\|\boldsymbol{\eta}\|_{2}^{2}.$
The term $\chi_{2}^{obs}$ is controlled using the same symmetrization and covering number arguments as in Theorem 4.1, giving
$\chi_{2}^{obs}\leq\frac{C_{3}mR_{obs}^{2}}{N}\log\mathcal{Q}\left(\frac{\delta}{C_{1}mR_{obs}},\mathcal{F}_{\mathrm{NN}}^{obs},\|\cdot\|_{L^{\infty,\infty}}\right)+3C_{1}\delta.$
Using the covering number estimate for neural networks and choosing $\delta=\varepsilon$, we obtain
$\chi_{2}^{obs}\leq\mathcal{O}\Big(D_{y}\log^{2}D_{y}\,\varepsilon^{-n}\log^{4}\varepsilon^{-1}\Big),$
and if $\mathbf{H}$ is learned, the hypothesis class is augmented by a linear operator in $\mathbb{R}^{m\times n}$, whose covering number contributes an additional $\mathcal{O}(mn\log\delta^{-1})$ term, affecting only the constants.
Combining the bounds yields
$\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{y})}\|\widehat{\mathscr{E}}_{obs}(\mathbf{y})-\mathbf{H}\,\mathscr{E}(\mathbf{x})\|_{2}^{2}\leq C_{0}^{obs}D_{y}^{2}\log^{2}D_{y}N^{-\frac{2}{n+2}}\log^{4}N+C_{1}^{obs}\,\mathbb{E}\|\tilde{\mathbf{\eta}}\|_{2}^{2},$
for some constant $C_{0}^{obs}>0$, which completes the proof.
Appendix C Algorithmic Details of Baseline Methods
Algorithm 2 AE-EnKF
1:Encoder $\mathscr{E}$, decoder $\mathscr{D}$, initial ensemble $\{\widehat{\mathbf{x}}_{0}^{(j)}\}_{j=1}^{N_{e}}$
2:for $k=1,\dots,K$ do
3:  Forecast: $\widehat{\mathbf{x}}_{k|k-1}^{(j)}=\mathscr{D}(\mathscr{E}(\widehat{\mathbf{x}}_{k-1}^{(j)}))$
4:  Analysis: Update $\widehat{\mathbf{x}}_{k|k-1}^{(j)}$ using EnKF with observation $\mathbf{y}_{k}$ and observation operator $\mathcal{H}$
5:end for
Algorithm 3 DAE-EnKF
1:Encoder $\mathscr{E}$, decoder $\mathscr{D}$, latent model $\mathscr{A}$, initial ensemble $\{\widehat{\mathbf{x}}_{0}^{(j)}\}_{j=1}^{N_{e}}$
2:Encode: $\widehat{\mathbf{z}}_{0}^{(j)}=\mathscr{E}(\widehat{\mathbf{x}}_{0}^{(j)})$
3:for $k=1,\dots,K$ do
4:  Forecast: $\widehat{\mathbf{z}}_{k|k-1}^{(j)}=\mathscr{A}(\widehat{\mathbf{z}}_{k-1}^{(j)})$
5:  Analysis: Update $\widehat{\mathbf{z}}_{k|k-1}^{(j)}$ using EnKF with observation $\mathbf{y}_{k}$ and observation operator $\mathcal{H}\circ\mathscr{D}$
6:end for
7:Decode: $\widehat{\mathbf{x}}_{k}^{(j)}=\mathscr{D}(\widehat{\mathbf{z}}_{k}^{(j)})$